I came across this question and they solved it by using the bar model.
How do I solve this algebraically? Why do you add 6+10 to get 16?
Mr. Lee wants to buy 5 shirts. If he buys 3 shirts, he will have $6 left. If he gets all of them, he will be $10 short. How much money does Mr. Lee have?
2 units: 6+10=16
1 unit: 16/2=8
Mr. Lee has $30
Mr. Lee has money for 3 shirts and $6 extra
To buy two more shirts, Mr. Lee has to use $6 leftover and find another $10 he is short of.
Thanks in advance.

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52776)

.

In accordance with the context of the problem, all the shirts have the same price.


Let x be the price for one shirt.


Then Mr. Lee has, from one part of the condition,  3x + 6 dollars;

                  from the other part of the condition, he has  5x - 10 dollars.


It gives you THIS EQUATION


    3x + 6 = 5x - 10.


Simplify and find x


    6 + 10 = 5x - 3x

    16     = 2x

     x     = 16/2 = 8 dollars.


One shirt costs 8 dollars.


Hence, Mr Lee has  3x + 6 = 3*8 + 6 = 30 dollars.    ANSWER

Solved.



Answer by MathTherapy(10551)


I came across this question and they solved it by using the bar model.
How do I solve this algebraically? Why do you add 6+10 to get 16?
Mr. Lee wants to buy 5 shirts. If he buys 3 shirts, he will have $6 left. If he gets all of them, he will be $10 short. How much money does Mr. Lee have?
2 units: 6+10=16
1 unit: 16/2=8
Mr. Lee has $30
Mr. Lee has money for 3 shirts and $6 extra
To buy two more shirts, Mr. Lee has to use $6 leftover and find another $10 he is short of.
Thanks in advance.
I'm not familiar with the bar model and don't have time to familiarize myself with it.
Anyway, let the amount the person has, be A, and the cost of each shirt, S
Then if 3 shirts are purchased, and $6 is left, we get the following equation: A - 3S = 6 ------ eq (i)
If all 5 shirts are purchased, he's short $10, so we get the following equation: A - 5S = - 10 ------ eq (ii)
In order to ELIMINATE A and find the cost of 1 shirt, we simply SUBTRACT eq (ii) from eq (i), which will then cause you to ADD 6 and 10.
As a matter of fact, when you perform the subtraction, "A" is eliminated and you get: 2S = 16
You would go on from there to find the cost of each shirt, and then further to find A, or the amount of money Mr. Lee has!
And, as you already know, that's supposed to be $30.